### v2

@@ 3.3.2 Coarse-Scale Processor
Figure 6: Schematic of synchronized message passing layer operations given by Eqs. 8-11. For ease of visualization, 2D projections of P=1 elements are shown and level index superscripts are ignored. (a) Coarse query element $G_{c}^{q}$ (blue box) and its neighbors ${\cal N}(G_{c}^{q})$ (red boxes) for the 26-neighbor case. Coincident owner nodes for purposes of illustration are denoted by red markers. (b) Edge update process (Eq. 8) on an edge in query element. (c) Edge aggregation process (Eq. 9) for owner node in query element. (d) Synchronization of edge aggregates as described in Eq. 10, with ${\cal C}(j)$ indicated by dashed box. (e) Node update process (Eq. 11) shown for owner node in query element. Note that in practice, the operations (b)-(d) are batched/parallelized over all nodes and edges on all graphs in the element neighborhood in (a).
@@ 3.3.3 Graph Unpooling/Interpolation Layer
The goal of the unpooling stage is to lift the embedded coarse-scale representation (the node features in the latent query graph $\widetilde{G}_{c}^{q}$) to the corresponding fine-graph representation in the same element. Such a lifting operation is evaluated within the GNN forward pass and is required to complete the element-local super-resolution procedure.
@@ 4 Results
With the super-resolution strategy described in Sec. 3, the goal of this section is to demonstrate the method on the mesh-based TGV dataset. Alongside assessment of the effect of coarse-element neighborhood size ${\cal N}_{i}$, the SRGNN architecture offers a useful capability to disentangle the effect of coarse-scale versus fine-scale message passing on the super-resolution procedure through proper specification of the number of layers used in the respective processors (see Fig. 5 and the surrounding discussion). As such, the demonstrations below leverage the following two model configurations:
@@
All models are implemented using a combination of PyTorch [66] and PyTorch Geometric [67] libraries. MLPs contain two hidden layers with ELU activations [68], layer normalization [69], and residual connections. Hidden node and edge feature dimensionalities are set to 128. Each model was trained for 100 epochs, with the training dataset at each Reynolds number consisting of 139968 query elements (coming from three separate instantaneous mesh-based snapshots), with 10% of all samples set aside during training for validation purposes. The Adam optimizer [70] was used to train all models with initial learning rate set to $10^{-4}$. A plateau-based learning rate scheduler (the ReduceLROnPleateau function) was used to dynamically adjust learning rates to mitigate effects of stagnation and increases in the validation loss. Models were trained using 8 Nvidia A100 GPUs (batch size of 4 per GPU) using two nodes of the Polaris supercomputer at the Argonne Leadership Computing Facility (ALCF). All results below are reported using predictions that have been unscaled/re-dimensionalized with respect to element-local statistics.
Below, GNN predictions are juxtaposed with an interpolation operation commonly used to initialize fine-scale flow-fields in multigrid cycles in the spectral element solution procedure. This is referred to as ”spectral element (SE) interpolation”, the methodology of which is provided in Ref. [61]. Although the SE interpolation is not explicitly designed for turbulent flow reconstruction, and is therefore not expected to compete with GNN-based reconstructions, it constitutes a useful reference point due to both its element-local nature (the SE interpolation can be interpreted as a linear prolongation operator acting on the element-local velocity fields, resembling the localized scope of the GNN operation) and the fact that it is already implemented in the NekRS flow solver. As such, reference comparisons to SE interpolation outputs will be made throughout this section.
4.1 Effects of GNN Configuration and Neighborhood Size
With regard to super-resolution accuracy, emphasis here is placed on comparison between the above described Model 1 and Model 2 configurations from the perspectives of instantaneous energy spectrum (global) analysis, element-local error analysis, and qualitative visualizations of super-resolved flow-fields. Alongside the distinction between Model 1 and Model 2 performance, discussion is centralized around the influence of coarse element neighborhood sizes on model prediction accuracy. Additionally, in this section, GNN super-resolution is assessed on snapshots unseen during the training stage in a fixed Reynolds number regime. In other words, models are trained and tested at the same Reynolds numbers to isolate capability of temporal generalization – assessment of Reynolds number extrapolations is provided in Sec. 4.2.
@@ 4.1.1 Global Error Analysis
To provide a global evaluation of the impact of model configuration, element neighborhood size, and Reynolds number on super-resolution accuracy, objective functions (in the form of average element-local mean-squared errors, as per Eq. 15) for the respective models evaluated on the unseen test set snapshots are shown in Fig. 7(a) and (b) for models trained (and evaluated on) Re=1600 and 3200 datasets, respectively. The figure reveals an increase in test set errors from Re=1600 to 3200 by a factor of roughly 2 in all model configurations, directly proportional to the increase in Reynolds number. This increase in error can be attributed to the significant added complexity in the Re=3200 dataset in terms of fine-scale turbulence structures relative to the Re=1600 case (see Fig. 2). The implication is that for a fixed model architecture (a fixed number of message passing layers), the super-resolution procedure encounters greater difficulty in direct proportion to the Reynolds number, a quality consistent with physical expectations.
At Re=1600, errors in each snapshot drop noticeably in Model 1 when moving from 0 to 6 neighbors, whereas in the Model 2 setting, less dependency is seen on the number of neighbors to the overall error. Physically, this implies that at Re=1600, the purely coarse-scale model (Model 1) relies on information from surrounding elements to achieve comparable super-resolution accuracy in a global sense to the multiscale model (Model 2). Interestingly, the same is not true for the Re=3200 models. In Fig. 7(b), errors are fairly insensitive to the neighborhood size, with some cases exhibiting inferior performance at 26 neighbors (and conversely, superior performance at 0 neighbors). As a result, at Re=3200, coarse-scale information in neighboring elements around the query element is hypothesized to be less helpful in reconstructing the flow-fields on the P=7 mesh. Although this phenomenon requires further investigation from the physical perspective, this may also imply that much of the ”retrievable” information content provided by the model form (the message passing operations) is contained within the query element at Re=3200. A more detailed analysis of errors (both coarse- and fine-scales) is ultimately needed, however, to understand the sources of these errors; such analysis is provided below.
@@
Figure 7: Query element MSE averaged over all elements in one-shot super-resolution task in test set snapshots for (a) models trained and evaluated using Re=1600 data and (b) models trained and evaluated using Re=3200 data. Each bar represents a separately trained GNN: errors shown for models using 0, 6, and 26 coarse element neighbors in both Model 1 (coarse-scale) and Model 2 (multiscale) GNN configurations.
To supplement the global MSE metric in Fig. 7, instantaneous energy spectra and corresponding errors in spectra predictions relative to the P=7 targets are shown in Fig. 8 and 9 for Re=1600 and 3200, respectively. Predictions are shown for two snapshots in each case: one in the training distribution (Fig. 8(a), Fig. 9(a)), and another in a time-extrapolated test regime (Fig. 8(b), Fig. 9(b)). In each scenario, results from GNNs trained using increasingly larger coarse element neighborhood sizes (0, 6, and 26) around the query element are shown.
@@
Assessment of model evaluations on the training snapshots in Fig. 8(a) and 9(a) serve as validation for the training procedure. Immediately apparent in both cases is that the GNN models are substantially more effective in recovering the target spectrum than the SE interpolation, which again is expected. It should be noted that the SE interpolation operator does recover some fine-scale information – it eliminates the accumulation of energy in the coarse P=1 solution in the inertial range and transfers this energy in a dispersed fashion into the high-wavenumber regime.
At Re=1600, both Model 1 and Model 2 spectrum curves here appear to be indistinguishable from the target P=7 spectrum on the training snapshot (Fig. 8(a)). However, through examination of corresponding relative errors, discrepancies between Model 1 (coarse-scale) and Model 2 (multiscale) performance can be seen for the training set. Specifically, the multiscale operations – through inclusion of message passing layers in the fine-scale processor – notably provide lower errors at the highest wavenumbers (without considering frequencies beyond the upper Nyquist limit), a result that is consistent with inclusion of fine-scale operations in the model form. On the training snapshot, the effect of neighborhood size is less pronounced: in Model 1, there is some error drop at the lower Nyquist limit (vertical blue line) going from 0 to 6 neighbors, and in Model 2, there is error drop near $k=100$ when moving from 6 to 26 neighbors. As will be described further below, the effects of neighborhood size – although less apparent in training results – come into play when generalizing models to time-extrapolated regimes. On the other hand, at Re=3200 (Fig. 9(a)), discrepancies between Model 1 and Model 2 are much more apparent in the training snapshot, with significantly higher relative errors observed in Model 1 (coarse-scale) on the training snapshot at the high wavenumbers. Unlike the Re=1600 counterpart in Fig. 8(a), there is less noticeable dependency on the neighborhood size at Re=3200; this overall insensitivity to neighborhood size for the Re=3200 case is consistent with the MSE findings presented for the testing snapshots in Fig. 7.
@@
Figure 8: (a) Top row: Instantaneous energy spectra for training set snapshot (t=10) at Re=1600. Curves shown for DNS target (black), coarse input (blue), spectral interpolation (gray), Model 1 (coarse-scale GNN, dashed-red), and Model 2 (multiscale GNN, dashed-green). Bottom row: Corresponding relative errors with respect to DNS spectrum versus wavenumber. (b) Same as (a), but for a time-extrapolated snapshot at t=10.5 also at Re=1600. As indicated by titles, plots from left to right correspond to increasing number of coarse element neighbors used in predictions (see Fig. 4).
@@
Figure 9: (a) Top row: Instantaneous energy spectra for training set snapshot (t=10) at Re=3200. Curves shown for DNS target (black), coarse input (blue), spectral interpolation (gray), Model 1 (coarse-scale GNN, dashed-red), and Model 2 (multiscale GNN, dashed-green). Bottom row: Corresponding relative errors with respect to DNS spectrum versus wavenumber. (b) Same as (a), but for a time-extrapolated snapshot at t=10.5 also at Re=1600. As indicated by titles, plots from left to right correspond to increasing number of coarse element neighbors used in predictions (see Fig. 4).
@@
Super-resolved spectra for test snapshots are shown in Fig. 8(b) and 9(b) for Re=1600 and 3200, respectively; these figures illustrate the generalization capability of the models in a time-extrapolated setting. At Re=1600, inspection of the spectrum curves (top row of Fig. 8(b)) shows how all models retain the ability to recover energy at the high-wavenumbers, and patterns in the low-wavenumber regimes are qualitatively similar to training set results. However, in the high-wavenumber region (particularly near $k=100$), the effect of slight over-predictions of energy content relative to the target is present at all neighborhood sizes. The relative error curves (bottom row) highlight key differences between training set predictions in two facets, which are less apparent on visual inspection of spectrum curves: although errors are lower than the SE interpolation for both Model 1 and Model 2 predictions, increases to relative errors appear in low-wavenumber regimes when compared to training set counterparts. These errors are non-monotonic with respect to $k$, with characteristic spikes near both low- and high-resolution Nyquist limits ($k=36$ and $k=100$, respectively).
@@
In the Re=3200 test set predictions (Fig. 9(b)), two key trends are apparent: (1) test set performance is nearly identical between Models 1 and 2 at this higher Reynolds number with respect to global energy spectrum recover, despite the observed training discrepancies shown in Fig. 9(a), and (2) for both models, the level of energy overshoot is smaller in the high wavenumber regime (near k=100) than in Re=1600. This reduction in overshoot at high wavenumbers in Re=3200 is balanced by an energy undershoot near the P=1 Nyquist limit at k=36. It is this undershoot at the relatively larger scales that leads to the overall higher global errors observed in Fig. 7 for Re=3200. Additionally, the relative error curves in Re=3200 are qualitatively very similar to those seen at Re=1600, which indicate the same characteristic buildup to a peak in the lower wavenumber range and a subsequent dip in the higher wavenumber range – the difference is that these features are horizontally shifted (i.e., the peak and dip in error occur at respectively higher wavenumbers), an effect of evaluating the same architecture at a higher Reynolds number. Overall, at the higher Reynolds number of Re=3200, the error trends between Model 1 and 2 configurations are quite similar, leading to the useful conclusion that in some settings, a model comprising purely coarse-scale learnable operations can be used to reconstruct fine-scale quantities of interest at a level comparable with a multiscale model.
The spectrum errors also provide insights on the sensitivities of models to the coarse element neighborhood sizes. For example, at Re=1600 (Fig. 8(b)), when increasing the neighborhood size beyond 0, the peak of relative error in the low-wavenumber regime counterintuitively increases below the coarse-scale Nyquist limit (as monitored by the first black circle outline in Fig. 8). Further, errors decrease in the high wavenumber regime (monitored by second black circle) with an increase to the neighborhood size. Minimum errors at the highest wavenumbers are lowest in Model 2 configurations at the larger neighborhood sizes, illuminating the advantages of including fine-scale message passing operations in the architecture. The Re=3200 spectra (Fig. 9(b)), on the other hand, show minimal sensitivity to neighborhood size, a quality that again mimics the global MSE trends shown in Fig. 7. Additionally, a strange trend is found in the errors for the 6-neighbor models at both Reynolds numbers, in that a noticeable spike is present at the lowest $k=1$ frequency – this may be due to the fact that the size-6 neighborhood does not completely surround the query element, introducing inconsistencies in message propagation throughout the neighborhood elements (this error disappears in the Model 2 configuration with 26 neighbors).
@@
The effect of neighborhood size on spectrum reconstruction is more accessible in Fig. 10, which plots relative error statistics versus number of coarse element neighbors in two wavenumber bins. These bins are outlined in Fig. 10(a), and are extracted based on the P=1 and P=7 Nyquist wavenumbers shown in the vertical lines; the first bin is in the low-wavenumber range where $k=[1,36]$ (Fig. 10(b)), and the other is in the high-wavenumber range where $k=[37,144]$ (Fig. 10(c)). Within these wavenumber ranges, average, maximum, and minimum observed values of spectrum relative errors from Model 1 and Model 2 predictions on the test set snapshots are shown.
@@
Figure 10: (a) Reference plot of spectrum relative error versus wavenumber (reproduced from Fig. 8). Low and high-wavenumber ranges indicated on top of plot. (b) Statistics of spectrum relative errors conditioned on low wavenumber range versus number of coarse element neighbors used in GNN predictions, with Re=1600 on top and Re=3200 on bottom. (c) Same as (b), but conditioned on high wavenumber range. In (b) and (c), mean is solid line, maximum is dot-dashed, and minimum is dashed, with Model 1 values in black circles and Model 2 in blue squares.
4.1.2 Local Error Analysis
Although assessment of performance from a global error analysis perspective is useful, additional physical insights related to model performance can be gained through error analysis at a more local level. In this vein, Fig. 11 provides a comparison of Model 1 and 2 element-local super-resolution error (SE interpolation errors are overall higher than the GNNs, and are therefore not shown in Fig. 11 for clarity). To expose underlying relationships between coarse-scale velocity field statistics and GNN prediction errors, the figure provides velocity standard deviations in the input query element (y-axis, a coarse scale statistic) versus super-resolution prediction errors (x-axis, a fine scale statistic) at the element-local level, for both Re=1600 and Re=3200 test data.
A more qualitative assessment of the mesh-based super-resolution procedure is provided in Fig. 12(a) and (b), which shows prediction visualizations on a testing set snapshot at Re=1600 and 3200 respectively. The figure juxtaposes Model 1 and Model 2 super-resolved fields with the SE interpolation, while providing reference visualizations for target (P=7) and coarse input (P=1) fields. The voricity contour visualizations are qualitative proxies for the turbulence intensity observed in the flow; comparison of these contours in both target P=7 and input P=1 flow-fields uncovers the inherent challenge in the super-resolution procedure, in that practically all of the turbulence is eliminated through the action of the coarse projection. Given the collection of input P=1 element graphs and their respective neighborhoods, the vorticity contour visualizations highlight the ability of GNNs to recover much of the target fine-scale structures at both tested Reynolds numbers, particularly when comparing with a conventional SE interpolation procedure.
Although the GNNs successfully generate fine-scale flow structures, closer inspection of the super-resolved fields (best visualized in the vorticity isosurfaces near the domain boundaries in the top row plots in Fig. 12) reveals non-physical noisy artifacts in the predictions. At Re=1600, lower levels of noise are observed in the Model 2 (multiscale) flowfields as compared to Model 1 (coarse-scale), with comparable levels of noise observed in both models at Re=3200. This mirrors the global trends in relative spectrum errors discussed in Sec. 4.1.1: the effect of these high-frequency artifacts is consistent with the presence of errors in the energy spectrum (including overshoots) observed in Figs. 8 and 9 in the high-wavenumber regimes.

### v3

@@ 3.3.2 Coarse-Scale Processor
Figure 7: Schematic of synchronized message passing layer operations given by Eqs. 5-8. For ease of visualization, 2D projections of P=1 elements are shown and level index superscripts are ignored. (a) Coarse query element $G_{c}^{q}$ (blue box) and its neighbors ${\cal N}(G_{c}^{q})$ (red boxes) for the 26-neighbor case. Coincident owner nodes for purposes of illustration are denoted by red markers. (b) Edge update process (Eq. 5) on an edge in query element. (c) Edge aggregation process (Eq. 6) for owner node in query element. (d) Synchronization of edge aggregates as described in Eq. 7, with ${\cal C}(j)$ indicated by dashed box. (e) Node update process (Eq. 8) shown for owner node in query element. Note that in practice, the operations (b)-(d) are batched/parallelized over all nodes and edges on all graphs in the element neighborhood in (a).
@@ 3.3.3 Graph Unpooling/Interpolation Layer
The goal of the unpooling stage is to lift the embedded coarse-scale representation (the node features in the latent query graph $\widetilde{G}_{c}^{q}$) to the corresponding fine-graph representation in the same element. Such a lifting operation is evaluated within the GNN forward pass and is required to complete the element-local super-resolution procedure. Note that an unpooling layer is also used outside of the core architecture to facilitate a residual-based GNN prediction (refer to Fig. 6) – this distinction is explained in D
@@ 4 Results
With the super-resolution strategy described in Sec. 3, the goal of this section is to demonstrate the method in various contexts. This is done in the following sections. First, in Sec. 4.1 and Sec. 4.2, the GNN performance is independently analyzed on the TGV and BFS datasets, respectively, focusing on model evaluations on unseen snapshots at fixed Reynolds numbers and geometry configurations (i.e., models are trained and tested at the same Reynolds numbers in these sections to isolate capability of temporal generalization). Then, in Sec. 4.3, assessment of models extrapolated to out-of-distribution Reynolds numbers is performed using the TGV configuration. Lastly, in Sec. 4.4, assessment of models in a geometry extrapolation context is performed using the cavity case.
Implementation details: All models are implemented using a combination of PyTorch [74] and PyTorch Geometric [75] libraries. MLPs contain two hidden layers with ELU activations [76], layer normalization [77], and residual connections. Hidden node and edge feature dimensionalities are set to 128. Each model was trained for 100 epochs, with 10% of all samples in respective datasets set aside during training for validation purposes. The Adam optimizer [78] was used to train all models with initial learning rate set to $10^{-4}$. A plateau-based learning rate scheduler (the ReduceLROnPleateau function) was used to dynamically adjust learning rates to mitigate effects of stagnation and increases in the validation loss. Models were trained using 8 Nvidia A100 GPUs (batch size of 4 per GPU) using two nodes of the Polaris supercomputer at the Argonne Leadership Computing Facility (ALCF). All results below are reported using predictions that have been unscaled/re-dimensionalized with respect to element-local statistics.
Spectral element interpolation: GNN predictions are juxtaposed with an interpolation operation typically used to initialize fine-scale flow-fields in multigrid cycles in the spectral element solution procedure. This is referred to as ”spectral element (SE) interpolation”, the methodology of which is provided in Ref. [54]. Although the SE interpolation is not explicitly designed for turbulent flow reconstruction, and is therefore not expected to compete with GNN-based reconstructions, it constitutes a useful reference point due to both its element-local nature (the SE interpolation can be interpreted as a linear prolongation operator acting on the element-local velocity fields, resembling the localized scope of the GNN operation) and the fact that it is already implemented in the NekRS flow solver. As such, reference comparisons to SE interpolation outputs will be made throughout this section.
Alongside assessment of the effect of coarse-element neighborhood size ${\cal N}_{i}$, the SRGNN architecture offers a useful capability to disentangle the effect of coarse-scale versus fine-scale message passing on the super-resolution procedure through proper specification of the number of layers used in the respective processors (see Fig. 6 and the surrounding discussion). As such, the TGV demonstrations leverage the following two model configurations:
@@
The total number of message passing layers in both Model 1 and 2 was set to 12 based on empirical tests that demonstrated good model performance up to this value. Additionally, at this number of message passing layers, and using the message passing variant described in Sec. 3.3, over-smoothing effects are not observed [40]. Training costs, in terms of total epoch times and model throughput extracted from TGV datasets, are shown in Table 1 as a reference.
GNN Configuration: Coarse element neighbors: Model 1 (Coarse-scale) Model 2 (Multiscale)
0 Nbrs 6 Nbrs 26 Nbrs 0 Nbrs 6 Nbrs 26 Nbrs
Training time, 1 epoch [sec] 128.2 s 144.7 s 151 s 141.9 s 150.4 s 155.2 s
Training throughput [elements/sec] 982.6 870.6 834.3 887.8 837.6 811.7
Table 1: GNN training time and throughput (in terms of elements processed per second during training). Data collected from Re=3200 TGV training runs, using 2 Polaris nodes (8 total Nvidia A100 GPUs). Total number of training elements was 125973 (this comes from roughly 90% of the total 139968 elements in the TGV dataset, with the remaining elements used as a validation set).
4.1 Demonstration on Taylor-Green Vortex (TGV)
With regard to super-resolution accuracy, emphasis here is placed on comparison between the above described Model 1 and Model 2 configurations from the perspectives of instantaneous energy spectrum (global) analysis and qualitative visualizations of super-resolved flow-fields. Alongside the distinction between Model 1 and Model 2 performance, discussion is centralized around the influence of coarse element neighborhood sizes on model prediction accuracy. Some additional analysis of element-local super-resolution error is provided in B.
@@ 4.1.1 Global Error Analysis
Element-Local Mean-Squared Errors: To provide a global evaluation of the impact of model configuration, element neighborhood size, and Reynolds number on super-resolution accuracy, objective functions (in the form of average element-local mean-squared errors, as per Eq. 12) for the respective models evaluated on the unseen test set snapshots are shown in Fig. 8(a) and (b) for models trained (and evaluated on) Re=1600 and 3200 datasets, respectively. The figure reveals an increase in test set errors from Re=1600 to 3200 by a factor of roughly 2 in all model configurations, directly proportional to the increase in Reynolds number. This increase in error can be attributed to the significant added complexity in the Re=3200 dataset in terms of fine-scale turbulence structures relative to the Re=1600 case (see A). The implication is that for a fixed model architecture (a fixed number of message passing layers), the super-resolution procedure encounters greater difficulty in direct proportion to the Reynolds number, a quality consistent with physical expectations. Additionally, at Re=1600, errors in each snapshot drop noticeably in Model 1 when moving from 0 to 6 neighbors, whereas in the Model 2 setting, less dependency is seen on the number of neighbors to the overall error.
@@
Figure 8: Query element MSE averaged over all elements in one-shot super-resolution task in test set snapshots for (a) TGV models trained and evaluated using Re=1600 data and (b) TGV models trained and evaluated using Re=3200 data. Each bar represents a separately trained GNN: errors shown for models using 0, 6, and 26 coarse element neighbors in both Model 1 (coarse-scale) and Model 2 (multiscale) GNN configurations.
@@
Energy Spectra: To supplement the global MSE metric in Fig. 8, instantaneous energy spectra and corresponding errors in spectra predictions relative to the P=7 targets are shown in Fig. 9 and 10 for Re=1600 and 3200, respectively. Assessment of model evaluations on the training snapshots in Fig. 9(a) and 10(a) serve as validation for the training procedure. Immediately apparent in both cases is that the GNN models are substantially more effective in recovering the target spectrum than the SE interpolation, which again is expected. It should be noted that the SE interpolation operator does recover some fine-scale information – it eliminates the accumulation of energy in the coarse P=1 solution in the inertial range and transfers this energy in a dispersed fashion into the high-wavenumber regime.
@@
Figure 9: (a) Top row: Instantaneous TGV energy spectra for training set snapshot (t=10) at Re=1600. Curves shown for DNS target (black), coarse input (blue), spectral interpolation (gray), Model 1 (coarse-scale GNN, dashed-red), and Model 2 (multiscale GNN, dashed-green). Bottom row: Corresponding relative errors with respect to DNS spectrum versus wavenumber. (b) Same as (a), but for a time-extrapolated snapshot at t=10.5 also at Re=1600. As indicated by titles, plots from left to right correspond to increasing number of coarse element neighbors used in predictions (see Fig. 5).
@@
Figure 10: (a) Top row: Instantaneous TGV energy spectra for training set snapshot (t=10) at Re=3200. Curves shown for DNS target (black), coarse input (blue), spectral interpolation (gray), Model 1 (coarse-scale GNN, dashed-red), and Model 2 (multiscale GNN, dashed-green). Bottom row: Corresponding relative errors with respect to DNS spectrum versus wavenumber. (b) Same as (a), but for a time-extrapolated snapshot at t=10.5 also at Re=1600. As indicated by titles, plots from left to right correspond to increasing number of coarse element neighbors used in predictions (see Fig. 5).
@@
Super-resolved spectra for test snapshots are shown in Fig. 9(b) and 10(b) for Re=1600 and 3200, respectively; these figures illustrate the generalization capability of the models in a time-extrapolated setting. At Re=1600, inspection of the spectrum curves (top row of Fig. 9(b)) shows how all models retain the ability to recover energy at the high-wavenumbers, and patterns in the low-wavenumber regimes are qualitatively similar to training set results. However, in the high-wavenumber region (particularly near $k=100$), the effect of slight over-predictions of energy content relative to the target is present at all neighborhood sizes. The relative error curves (bottom row) highlight key differences between training set predictions in two facets, which are less apparent on visual inspection of spectrum curves: although errors are lower than the SE interpolation for both Model 1 and Model 2 predictions, increases to relative errors appear in low-wavenumber regimes when compared to training set counterparts. These errors are non-monotonic with respect to $k$, with characteristic spikes near both low- and high-resolution Nyquist limits ($k=36$ and $k=100$, respectively).
@@
In the Re=3200 test set predictions (Fig. 10(b)), two key trends are apparent: (1) test set performance is nearly identical between Models 1 and 2, despite the observed training discrepancies shown in Fig. 10(a), and (2) the level of energy overshoot is smaller in the high wavenumber regime (near k=100) than in Re=1600. This is balanced by an energy undershoot near the P=1 Nyquist limit at k=36. It is this undershoot at the relatively larger scales that leads to the overall higher global errors observed in Fig. 8 for Re=3200. Additionally, the relative error curves in Re=3200 are qualitatively very similar to those seen at Re=1600, which indicate the same characteristic buildup to a peak in the lower wavenumber range and a subsequent dip in the higher wavenumber range – the difference is that these features are horizontally shifted (i.e., the peak and dip in error occur at respectively higher wavenumbers), which is an effect of evaluating the same architecture at a higher Reynolds number. Overall, at the higher Reynolds number of Re=3200, the error trends between Model 1 and 2 configurations are quite similar, leading to the useful conclusion that in some settings, a model comprising purely coarse-scale learnable operations can be used to reconstruct fine-scale quantities of interest at a level comparable with a multiscale model.
4.1.2 Effect of Neighborhood Size
@@
The effect of neighborhood size on spectrum reconstruction is provided in Fig. 11, which plots relative error statistics versus number of coarse element neighbors in two wavenumber bins. These bins are outlined in Fig. 11(a), and are extracted based on the P=1 and P=7 Nyquist wavenumbers shown in the vertical lines; the first bin is in the low-wavenumber range where $k=[1,36]$ (Fig. 11(b)), and the other is in the high-wavenumber range where $k=[37,144]$ (Fig. 11(c)). Within these wavenumber ranges, average, maximum, and minimum observed values of spectrum relative errors from Model 1 and Model 2 predictions on the test set snapshots are shown.
Figure 11: (a) Reference plot of TGV spectrum relative error versus wavenumber (reproduced from Fig. 9). Low and high-wavenumber ranges indicated on top of plot. (b) Statistics of spectrum relative errors conditioned on low wavenumber range versus number of coarse element neighbors used in GNN predictions, with Re=1600 on top and Re=3200 on bottom. (c) Same as (b), but conditioned on high wavenumber range. In (b) and (c), mean is solid line, maximum is dot-dashed, and minimum is dashed, with Model 1 values in black circles and Model 2 in blue squares.
@@
4.1.3 Physical Space Visualizations
A more qualitative assessment of the mesh-based super-resolution procedure is provided in Fig. 12(a) and (b), which shows prediction visualizations on a testing set snapshot at Re=1600 and 3200 respectively. The figure juxtaposes Model 1 and Model 2 super-resolved fields with the SE interpolation, while providing reference visualizations for target (P=7) and coarse input (P=1) fields. The vorticity contour visualizations are qualitative proxies for the turbulence intensity observed in the flow; comparison of these contours in both target P=7 and input P=1 flow-fields uncovers the inherent challenge in the super-resolution procedure, in that practically all of the turbulence is eliminated through the action of the coarse projection. Given the collection of input P=1 element graphs and their respective neighborhoods, the vorticity contour visualizations highlight the ability of GNNs to recover much of the target fine-scale structures at both tested Reynolds numbers, particularly when comparing with a conventional SE interpolation procedure.
Figure 12: Instantaneous mesh-based flow-fields on time-extrapolated TGV snapshot (t=10.5s) for (a) Re=1600 and (b) Re=3200. From left-to-right: P=7 target solution (DNS), P=1 input coarse solution (projected DNS), spectral element interpolation, GNN in Model 1 configuration (coarse-scale) using 26 neighbors, GNN in Model 2 configuration (multiscale) using 26 neighbors. Top row shows z-component vorticity contours in 3D (blue and red indicate vorticity of 3 and -3 respectively), middle row shows z-component of vorticity in the X-Y plane at z=$\pi/2$, and bottom row shows velocity magnitude in the same 2D plane. Black and white circles in 2D visualizations are provided as visualization guidelines.
Although the GNNs successfully generate fine-scale flow structures, closer inspection of the super-resolved fields (best visualized in the vorticity isosurfaces near the domain boundaries in the top row plots in Fig. 12) reveals non-physical noisy artifacts in the predictions. At Re=1600, lower levels of noise are observed in the Model 2 (multiscale) flow-fields as compared to Model 1 (coarse-scale), with comparable levels of noise observed in both models at Re=3200. This mirrors the global trends in relative spectrum errors discussed in Sec. 4.1.1: the effect of these high-frequency artifacts is consistent with the presence of errors in the energy spectrum (including overshoots) observed in Figs. 9 and 10 in the high-wavenumber regimes.